The specific heat capacity for silver is 0.24 J/goC. Calculate the energy required to raise the temperature of 150.0 g Ag from 273 K to 298 K.





1
Expert's answer
2022-04-19T03:11:58-0400

Solution:

q = m × C × ΔT

q = m × C × (Tf − Ti),

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J °C−1 g−1)

Tf = final temperature (°C)

Ti = initial temperature (°C)


m = 150.0 g

C = 0.24 J °C−1 g−1

Tf = 298 K = 25°C

Ti = 273 K = 0°C


Therefore,

q = (150.0 g) × (0.24 J °C−1 g−1) × ( 25°C − 0°C) = 900 J

q = 900 J


Answer: 900 J of energy required